OpenStudy (anonymous):

can someone please help me solve equation 4w^2=4w+3 for x? I have the answers but I don't understand the process to get to them. Please help

OpenStudy (anonymous):

what are your answers? @CadenN

OpenStudy (anonymous):

w= -3/2 and w=-1/2 the question was supposed to say solve for w, sorry

OpenStudy (anonymous):

well those answers are right you just have two answers.. you need to choose which one is best, umm let me try solving

OpenStudy (anonymous):

no, they're both right

OpenStudy (anonymous):

i just don't understand how to get to them from the exuation given

OpenStudy (anonymous):

i can solve your equations but they come up with those two answers, do you want me to write out step by step?

OpenStudy (anonymous):

yes please that would be a big help @cutelilcigarette

OpenStudy (anonymous):

@CadenN 4w^2 = 4w + 3 subtract 4w + 3 from both sides 4w^2 - (4w + 3) = 4w + 3 - (4w + 3) 4w^2 - 4w - 3 = 0 factor left side of equation (2w + 1)(2w - 3) = 0 set factors equal to 0 2w + 1 = 0 or 2w - 3 = 0 and this is where the two answers come in w = -1/2 or w = 3/2

OpenStudy (anonymous):

thank you, do you by any chance know what to do if 'a' was not a factorable number like 4 or 9 but for example 5?